### v1

# DualMAR: Medical-Augmented Representation from Dual-Expertise Perspectives

Abstract
Electronic Health Records (EHR) has revolutionized healthcare data management and prediction in the field of AI and machine learning. Accurate predictions of diagnosis and medications significantly mitigate health risks and provide guidance for preventive care. However, EHR driven models often have limited scope on understanding medical-domain knowledge and mostly rely on simple-and-sole ontologies. In addition, due to the missing features and incomplete disease coverage of EHR, most studies only focus on basic analysis on conditions and medication. We propose DualMAR, a framework that enhances EHR prediction tasks through both individual observation data and public knowledge bases. First, we construct a bi-hierarchical Diagnosis Knowledge Graph (KG) using verified public clinical ontologies and augment this KG via Large Language Models (LLMs); Second, we design a new proxy-task learning on lab results in EHR for pretraining, which further enhance KG representation and patient embeddings. By retrieving radial and angular coordinates upon polar space, DualMAR enables accurate predictions based on rich hierarchical and semantic embeddings from KG. Experiments also demonstrate that DualMAR outperforms state-of-the-art models, validating its effectiveness in EHR prediction and KG integration in medical domains.
Keywords EHR Prediction  $\cdot$ Diagnosis KG  $\cdot$ Medical-domain knowledge  $\cdot$ LLM  $\cdot$ Pretraining  $\cdot$ Lab results
1 Introduction
The digitization of patient information such as EHR has transformed healthcare in terms of data storage, information retrieval, and computational pattern recognition. EHR-based prediction models, including those for diagnosis, heart failure, and readmission prediction tasks (Shickel et al., 2017), leverage EHR data to forecast patient outcomes. In addition to improving predictive accuracy, a major focus of current research is to build models that understand medical domains from both individual health records (Poulain and Beheshti, 2024) and public knowledge bases (Jiang et al., 2023).
Individual health records preserve hidden knowledge of medical concepts, such as disease complications. Many models use EHR data to uncover disease progression patterns and learn medical embeddings, which serve as local-domain priors for downstream tasks. Meanwhile, transformer-based models following G-BERT (Shang et al., 2019) rely on self-supervised tasks to fully utilize EHR data, including single admissions, enhancing training efficiency and robustness of models. However, these methods typically focus on condition, treatment, and medication concepts, often overlooking critical information like lab tests.
(a) Generating Augmented Bi-hierarchical KG.
(b) The Dual-Expertise Module for EHR Prediction.
Figure 1: Overview of the proposed DualMAR framework.
From another perspective, public medical knowledge have been used in machine learning to enhance understanding of medical concepts Choi et al. (2017); Ma et al. (2018). Recent studies still primarily focused on simple hierarchical ontologies, like ICD-9 hierarchy Lu et al. (2023), showing that even simple and sole relations can help models get state-of-the-art performance in predictive healthcare. However, KGs with more complex and diverse relations remain under-explored in EHR models. To the best of our knowledge, GraphCare Jiang et al. (2023) is one of the preliminary attempts to use large language models such as GPT-4 Achiam et al. (2023) to generate KGs for EHR predictions, outperforming most existing models. Experimental results demonstrate that KGs refine prediction through comprehensive understanding for medical concepts, making them a promising direction for augmenting EHR models, though some unsolved challenges remain:
  1. 1.
    Data Quality of KG: GraphCare suffers from data quality issues with its generated triples. First, triples generated by GPT in GraphCare can lead to hallucinations upon zero-shot learning. Additionally, the final KG used in GraphCare is constrained by its reliance on the CCS Wei et al. (2017) encoding system, limiting its scope (# diseases: 285) for more specific diagnostic tasks.
  2. 2.
    Integration of KG into EHR Models: First, GraphCare initializes each patient with a personalized KG for prediction, which imposes demands on GPU memory and limits scalability as KG sizes increase. Second, GraphCare primary focus on the semantic information of medical concepts leads to a partial understanding within the KG, which, combined with its lack of hierarchical features, can result in biased predictions for EHR tasks.
To address these challenges in integrating KGs into predictive models, we propose leveraging both medical domain knowledge from generated KGs and a novel pretraining proxy task in EHR, which enables models to take advantage of both “knowledge scholar” and “local expert”. Our proposed method, DualMAR, is a framework which not only enhance embedding by retrieving both contextual and hierarchical information from KGs but also further refine medical representation through EHR graphs and pretraining modules. The main contributions of this work are as follows:
  • “Knowledge Scholar”: We establish a diagnosis-driven KG across public medical databases and design a bi-hierarchical structure that organizes core domains based on phenotype and condition codes, further enriching it with verified textual information through in-context learning with LLMs. We then employ polar-space projection to retrieve embeddings that capture both semantic and hierarchical information from generated KG.
  • “Local Expert”: We employ a Graph Learning module using a disease-lab EHR graph to enhance embeddings of medical codes, and introduce a novel proxy task to augment patient hidden representations from a laboratory perspective, fully leveraging the entire EHR dataset.
  • “Dual-expertise Perspective”: An encoder-decoder architecture is adopted for integrating KG into the predictive model. Embeddings provided by the “Knowledge Scholar” serve as prior knowledge for the “Local Expert” who continually refines these representations. The “Local Expert” converts patient admission history into laboratory embeddings through the encoder, resulting in accurate predictions across various downstream tasks.
2 Related Work
Deep learning models have been extensively applied in EHR prediction to provide guidance for preventive care. Both RNN-based Choi et al. (2016a, b); Ma et al. (2017); Bai et al. (2018); Ma et al. (2020a) and CNN-based models Nguyen et al. (2017); Ma et al. (2020b) have been developed to capture temporal and spatial dependencies for health risk predictions. With the advent of graph neural networks (GNNs), GRAM Choi et al. (2017) and its successors Liu et al. (2020); Wu et al. (2021); Lu et al. (2023) incorporate graph learning to further enhance hidden representation for medical concepts. Meanwhile, previous work such as G-BERT Shang et al. (2019) takes advantage of transformers and their variants to fully utilize EHR data (i.e. single-admission patients) through various proxy tasks Prakash et al. (2021); Rasmy et al. (2021); Poulain and Beheshti (2024).
Predictive models that leverage graph and transformer architectures have achieved state-of-the-art performance in various EHR tasks, highlighting their potential to advance predictive healthcare with existing knowledge. This knowledge can be divided into local individual health data and public knowledge bases, as outlined below:
  • Individual Health Data: Prior work such as CGL Lu et al. (2021) has constructed EHR graphs from EHR data based on co-occurrence of medical concepts in admission records to capture hidden relations among diseases and improve health outcome predictions. Encoder-decoder models following G-BERT Shang et al. (2019), focus on customized pretraining tasks to enhance the understanding of medical patterns in EHR data. While individual knowledge aids in adapting models to different datasets, it relies on the size of samples and the quality of data.
  • Public Knowledge Bases: External knowledge is derived from public data sources without individual information. Researchers have explored simple ontologies like the ICD-9 hierarchy, Retrieval-Augmented methods like RAM-EHR Xu et al. (2024) for unstructured medical descriptions, and KGs for predictive models like GraphCare Jiang et al. (2023). However, these approaches often focus solely on either hierarchical or semantic information, leading to a narrow understanding of medical concepts and potential information loss, and how to effectively combine both information in models is necessary to enhance robustness of predictive healthcare.
In addition to leveraging medical-domain knowledge, incorporating both structured and unstructured features can enhance predictions. MiME Choi et al. (2018) and GCT Choi et al. (2020) use graph structures with lab results to augment representations, while CGL Lu et al. (2021) and MedGTX Park et al. (2022) integrate unstructured clinical notes to highlight the value of additional information. However, most approaches rely heavily on these multi-aspect features, limiting their applicability to data without those features. Our proposed framework integrates laboratory features as auxiliary data in graph construction and proxy tasks, enhancing predictions even when no desired feature is available.
3 Methodology
We present DualMAR, a framework designed for predictive healthcare that aims to achieve a comprehensive understanding of diagnosis-related entities from both external (public) and local (individual) knowledge bases. An overview of the proposed framework is shown in Figure 1(b).
3.1 General Notation
An EHR dataset $\mathcal{S}$ is a collection of temporal admission records of $N$ patients $\{\mathbf{p}_{1},\mathbf{p}_{2},...,\mathbf{p}_{N}\}$, and $\mathbf{p}_{u}$ is denoted as certain patient $u$. Such patient can be also represented as a sequence of $T_{u}$ admission records $(V_{1}^{u},V_{2}^{u},..,V_{T}^{u})\in\mathbf{p}_{u}$ in chronological order. We omit $u$ to identify different patients in following sections and explain our framework using single patient as example to avoid misunderstanding. We denote the entire set of medical concepts as $\mathcal{C}=\{c_{1},c_{2},...,c_{|\mathcal{C}|}\}$ where $|\mathcal{C}|$ is the vocabulary size of medical concepts. Each admission, like $t$-th admission $V_{t}\in(V_{1},V_{2},..,V_{T})$, contains a subset of $\mathcal{C}$. Note that, We consider diseases and lab results as medical concepts in proposed framework, and we denote diseases $\{d_{1},d_{2},...,d_{|\mathcal{D}|}\}$ and lab results $\{l_{1},l_{2},...,l_{|\mathcal{L}|}\}$ within vocabulary $\mathcal{D},\mathcal{L}\subset\mathcal{C}$.
3.2 Generate “Knowledge Scholar”
In general, the proposed model integrates external knowledge from professional medical datasets and LLM-retrieved KGs based on verified clinical text. Figure 1(a) illustrates how to augment structured KG with textual information.
(1) HKG Construction: Healthcare KGs (HKGs) are healthcare-specific KGs that span multiple datasets (Cui et al., 2023), offering a more comprehensive understanding than all-in-one ontologies like UMLS Bodenreider (2004). We focus on constructing HKGs that link diseases, drugs, and phenotypes to enable precise clinical predictions. The Human Phenotype Ontology (HPO) and the ICD-9 hierarchy serve as the foundation for our KG, with bi-hierarchical structure enhancing its robustness. Inspired by previous healthcare KGs (Su et al., 2023; Chandak et al., 2023; Ghorbanali et al., 2023), we also incorporate updated and verified sources like Drug Central, DrugBank, and SIDER to create a highly connected HKG. We denote the resulting healthcare KG, sourced from existing databases, as $\mathcal{G}_{M}$, containing $M$ tuples of head, relation, and tail. More details can be found in Appendix A.
(2) LLM-KG Generation: We extract textual features for specific diseases from widely-used knowledge bases and then employ a template with instruction, example, and prompt to harness LLMs’ in-context learning capabilities. For instance, when generating a KG for a diagnostic concept, the LLM generates triples like “[Heart Failure, CAUSED_BY, Hypertension]” based on the collected text. Iterative prompting and re-reading strategies are used to refine outputs. The aggregated results form an LLM-generated KG $\mathcal{G}_{N}$. Further details on knowledge sources and prompting are in Appendices A.2 and A.3 respectively.
$\mathcal{G}_{M}$ $=\bigcup\nolimits_{M}(\mathbf{h}^{(M)}_{m},\mathbf{r}^{(M)}_{m},\mathbf{t}^{(M)}_{m})$ (1)
$\mathcal{G}_{N}$ $=\bigcup\nolimits_{x,y}\bigcup\nolimits_{N}(\mathbf{h}^{(N)}_{n},\mathbf{r}^{(N)}_{n},\mathbf{t}^{(N)}_{n})$ (2)
$\mathcal{G}_{H}$ $=\mathcal{G}_{M}\cup\mathcal{G}_{N},\>\tilde{\mathcal{G}_{H}}=\textsc{Normalize}(\mathcal{G}_{H})$ (3)
Here, $x$ means the number of iteration for iterative running and and $y$ denotes the number of times the generated output is re-evaluated (re-reading) to refine the quality of triples.
3.2.1 Node and Edge Normalization.
Our approach for normalizing the combined KG $\mathcal{G}_{H}$ involves two key steps: (1) Uniform encoding system for entities: In addition to phenotype nodes sourced from HPO, we align disease and drug entities using the ICD-9-CM Organization et al. (1988) and ATC Nahler and Nahler (2009) coding systems, both of which are widely used in EHR data. A cross-referencing method is employed to convert $\mathcal{G}_{M}$ and $\mathcal{G}_{N}$ into a shared hierarchy. Given that drugs are not the primary focus, we use ATC-4 to simplify the numerous synergistic interactions in DrugBank. (2) Clustering duplicated or unmatched triples: The LLM-prompting process is run multiple times, which can generate duplicated nodes and edges. Moreover, cross-referencing files may not cover all concepts, and unmatched entities leave noise in KG. To address these problems, we calculate cosine similarity for nodes and edges based on their embeddings. Using a predefined threshold $\theta$, we group entities within categories such as “disease”, “drug”, “phenotype”, and “other”, and apply hierarchical clustering to reorganize edges.
Table 1: Statistical Comparison among KGs.
(* means results might fluctuate by different parameter settings.)
KG #Nodes #EdgeTypes #Triples
GraphCare: [Disease, Procedure, Drug, Other]
   - Ontology* 10,805 54 81,073
   - GPT-4 output* 4,599 752 31,325
   - Final KG* 12,284 785 104,460
PrimeKG: [Disease, Drug, Protein … 10 in total]
   - Public KG 129,375 30 4,050,249
   - Text Description - - -
Diagnosis KG: [Disease, Drug, Phenotype, Other]
   - Ontology 32,941 6 477,992
   - LLM output* 42,436 3,642 82,191
   - Final KG* 52,604 3,645 560,183
3.2.2 Advancement Analysis
To demonstrate the advancements of our proposed Diagnosis KG, we compare it with two recent health-related KGs: GraphCare Jiang et al. (2023) and PrimeKG Chandak et al. (2023). A statistical summary of KGs is presented in Table 1. The analysis highlights the following advantages of our KG:
  1. 1.
    Compared to GraphCare, our KG is larger in scale and built on the widely-used ICD-9-CM system, offering hierarchical and precise knowledge for predictive healthcare. Additionally, while GraphCare relies on zero-shot generation and subgraph sampling, our KG provides more stable and accurate results.
  2. 2.
    Compared to PrimeKG, our KG includes more edge types, primarily augmented by LLM-generated triples. The LLM reasoning process not only enhances connectivity within the KG but also converts unstructured text into structured triples, making information retrieval more efficient from single-modal data.
Although PrimeKG covers a broader range of medical concepts, it is more prone to data quality issues that can introduce bias into EHR predictions. For example, genes and proteins included in PrimeKG are seldom considered in physicians’ decision-making processes, particularly when it comes to diagnosis. In contrast, our diagnosis-driven KG contains numerous normalized triples with diverse edges, enhanced by LLMs’ in-context learning capabilities, making it highly reproducible and well-suited for various diagnosis-related healthcare predictions.
3.3 Polar-space KG Embedding
Our generated KG features a bi-hierarchical structure, with diagnostic and phenotype nodes organized across multiple levels. Traditional KG embedding methods often focus primarily on contextual information, but our framework gives equal importance to both contextual and hierarchical information. We employ a polar-space embedding Zhang et al. (2020) method upon the Euclidean space due to its superiority in efficiently capturing hierarchical structures compared to other Euclidean embedding. While non-Euclidean methods are also designed for hierarchical structures, they primarily focus on hierarchy alone. In contrast, polar-space embedding not only capture both hierarchical and contextual information simultaneously but also provide a more computationally efficient and interpretable framework, making it well-suited for integrating complex medical knowledge into predictive models. A comparison among KG embedding methods Bordes et al. (2013); Nickel and Kiela (2017); Sun et al. (2019); Fionda and Pirrò (2020) is provided in Appendix B.
To describe medical concepts from both contextual and hierarchical perspectives, we leverage two properties of the polar coordinate system: (1) Radial coordinate: map entities across different levels of the hierarchy; (2) Angular coordinate: map entities at the same level based on contextual information. We denote entities, including head $\mathbf{h}$ and tail $\mathbf{t}$, as $\mathbf{e}$, and relation embeddings as $\mathbf{r}$. Subscripts, such as $\mathbf{e}_{r}$ and $\mathbf{e}_{a}$, represent the modulus and phase components, respectively. Each entry of $\mathbf{h}_{r,a}$ or $\mathbf{t}_{r,a}$ corresponds to a radial or angular value, while each entry of $\mathbf{r}_{r,a}$ represents a scaling transformation between the corresponding head and tail. The radial and angular coordinates of the entity embeddings can then be formulated as follows:
$\mathbf{h}_{r}\circ\mathbf{r}_{r}=\mathbf{t}_{r},\>\ \mathbf{h}_{r},\mathbf{t}_{r},\mathbf{r}_{r}\in\mathbb{R}^{k}$ (4)
$(\mathbf{h}_{a}+\mathbf{r}_{a})\bmod 2\pi=\mathbf{t}_{a},\>\ \mathbf{h}_{a},\mathbf{t}_{a},\mathbf{r}_{a}\in[0,2\pi)^{k}$ (5)
According to the coordinates for triples, we can further calculate radial and angular distances as follows:
$d_{r}(\mathbf{h}_{r},\mathbf{t}_{r})$ $=\|\mathbf{h}_{r}\circ\mathbf{r}_{r}-\mathbf{t}_{r}\|_{2},$ (6)
$d_{a}(\mathbf{h}_{a},\mathbf{t}_{a})$ $=\|\sin((\mathbf{h}_{a}+\mathbf{r}_{a}-\mathbf{t}_{a})/2)\|_{1}$ (7)
where $\|\cdot\|_{1}$, $\|\cdot\|_{2}$ denote the $\ell_{1}$, $\ell_{2}$ norm. The final circular arc distance $d(\mathbf{h},\mathbf{t})$ can be consider as the weighted sum of both radial and angular distances. Note that, we can get $\mathbf{h}$ and $\mathbf{t}$ by concatenating both radial and angular vectors. Such embedding model then be pretrained individually through link prediction tasks before integrating into the predictive model, which can help us understand priors from KG into trainable embeddings. The regular negative sampling loss function $L$ is adopted for optimization:
$L=-\log\sigma(\gamma-d(\mathbf{h},\mathbf{t}))-\sum_{i=1}^{n}\log\sigma(d(\mathbf{h}^{\prime},\mathbf{t}^{\prime})-\gamma)$ (8)
where $\gamma$ and $\sigma$ are marginal value and sigmoid function respectively, and $({h^{\prime}},r,{t^{\prime}})$ is a negative triple. Therefore, we can get diagnosis embedding matrix $\mathbf{E_{\mathcal{D}}}$ for condition codes, which is the initial nodes features for “Local Expert”.
3.4 Continual Enhancement from “Local Expert”
The continual enhancement process unfolds in two steps: graph learning and proxy-task learning, which together refine the priors from the “Knowledge Scholar” by fully leveraging EHR data. The proposed encoder-decoder framework not only emphasizes the comprehensive understanding of medical patterns but also seamlessly integrates local knowledge into the learning process, further capturing critical medical patterns for precise diagnosis.
3.4.1 Encoder Module.
The backbone of the encoder module contains graph learning and attention, which adjusts code-level embeddings and progressively refines them into comprehensive patient representations.
We construct a disease-complication graph $\mathcal{G}$ including both condition and laboratory codes. Compared to regular complication graphs involving only diseases, we have broader scope to explain complication in terms of shared abnormal lab tests in the same admission. To represent the complications between two diseases via patient admissions, we add bidirectional edges $(c_{i},c_{j})$, $(c_{i},c_{m})$, and $(c_{j},c_{m})$ in graph $\mathcal{G}$, where $c_{i},c_{j}\in\mathcal{D}$ are diagnostic codes and $c_{m}\in\mathcal{L}$ is a laboratory code. Therefore, we then compute co-occurrence matrix $\mathbf{B}\in\mathbb{N}^{|\mathcal{C}|\times|\mathcal{C}|}$ and adjacent matrix $\mathbf{A}=(1-\varphi)\mathbf{B}+\varphi\mathbf{I}\in\mathbb{R}^{|\mathcal{C}|\times|\mathcal{C}|}$ considering self loop based on EHR graph $\mathcal{G}$, where $\mathbf{I}$ is an identity matrix.
A GNN with $L$ layers is adopted by EHR graph $\mathcal{G}$ and node features $\mathbf{E}$, where we initialize diseases by priors $\mathbf{E_{\mathcal{D}}}$ from KG. Then, the hidden representation $\mathbf{H}^{(l)}$ can be calculated by input $\mathbf{H}^{(l-1)}$ through the $l$-th Convolution layer: After the last GNN layer, we obtain the hidden representation of diseases $\mathbf{X}=\textsc{GNN}(\mathbf{A},\mathbf{E})=\mathbf{H}^{(L)}$, which can be further refined through the bi-attention mechanism across codes and admissions for patient embedding $\mathbf{p}$:
$\mathbf{v}_{\tau}=\sum_{i=1}^{n}\alpha_{\tau}^{i}\mathbf{x}_{i},\quad\mathbf{p}=\sum_{\tau=1}^{T}\beta_{\tau}\tilde{\mathbf{v}}_{\tau},$ (9)
where $\alpha_{\tau}^{i}$ and $\beta_{\tau}$ are attention scores computed as follows:
$\mathbf{z}_{i}=\tanh(\mathbf{W}_{c}\mathbf{x}_{i}),\quad\mathbf{r}_{\tau}=\tanh(\mathbf{W}_{v}\sigma(\mathbf{W}_{u}\mathbf{v}_{\tau}))$ (10)
$\alpha_{\tau}^{i}=\frac{\exp(\mathbf{z}_{i})}{\sum_{j=1}^{n}\exp(\mathbf{z}_{j})},\quad\beta_{\tau}=\frac{\exp(\mathbf{r}_{\tau})}{\sum_{\tau=1}^{T}\exp(\mathbf{r}_{\tau})}$ (11)
$\mathbf{W}_{c}$, $\mathbf{W}_{u}$, and $\mathbf{W}_{v}$ are weighted matrices. The attention score $\alpha_{\tau}$ measures the distribution of medical codes of admission, and $\beta$ measures the distribution across admissions. Note that, we project each admission $\mathbf{v}_{\tau}$ to $\tilde{\mathbf{v}}_{\tau}$ for fitting the patient dimension through a single layer with $\mathbf{W}_{u}$. Detailed explanation about attention is provided in Appendix C.
The proposed encoder module ensures that both the specificity of individual medical encounters and the comprehensive history of a patient’s health are reflected in the final patient embedding $\mathbf{p}$, laying a robust foundation for the subsequent self-supervised learning phase.
3.4.2 Proxy-Task Learning.
Given patient embedding $\mathbf{p}$, we introduce a novel proxy-task learning paradigm. Some predictive models in EHR focus on customized proxy tasks to utilize single admission which is discarded in training process. However, designed labels of proxy tasks should not only reflect historical pattern of previous admissions, but also closely connect to downstream labels for learned patient embeddings $\mathbf{p}$. Therefore, we harness the insight of lab results in our learning model, which is unexplored as auxiliary features within pretraining module. By pretraining the model on laboratory results, we aim to mimic the diagnostic reasoning process of real physicians, who often rely on lab data for accurate assessments.
We denotes lab results as $\mathcal{L}$, and the designed proxy task makes predictive model inherently understands intricate relations between abnormal lab tests and health conditions. Considering patterns always change across different lab tests, we further categorize lab results into distinct groups, which can fully cover involved test items in EHR data. Here we take lab tests in MIMIC as an example:
  1. 1.
    Hematology (# items: 417): Analyze blood components like red and white blood cells to monitor conditions.
  2. 2.
    Chemistry (# items: 286) Evaluate chemical factors in blood and provide information about organ’s function.
  3. 3.
    Blood Gas (# items: 35) Measure the levels of gas in blood to assess respiratory and metabolic functions.
Hence, we categorize lab tests as $\mathcal{L}_{1}$, $\mathcal{L}_{2}$, and $\mathcal{L}_{3}$ in order, and predict lab results $\hat{y}_{i,j}=\text{Decoder}(\mathbf{p})$ based on both single-admission and multi-admission data in training set, where $y_{i,j}=1$ means lab code $l_{j}$ shows abnormal results, and a multi-layer perceptron (MLP) is used as decoder for learning patterns in different lab categories. We only focus on abnormal lab tests across admissions, so the ground truth $[l_{1},l_{2},...,l_{|\mathcal{L}|}]$ are labeled as multi-hot vectors. To ensure the lab embeddings are well-aligned with each category, the proxy task is designed to be completed in two steps:
  • Joint Training. We can get preliminary trained parameters in encoder $\Theta_{e}$ and optimize prediction through the average binary cross entropy (BCE) loss functions across three categories, where we can adjust patient embedding $\tilde{\mathbf{p}}$ through updated parameters.
    $L_{j}=\frac{1}{|\mathcal{Y}|}\sum_{i=1}^{|\mathcal{Y}|}\text{BCE}(\mathbf{\hat{y}}_{i},\mathbf{y}_{i}),\>\mathcal{Y}=\{\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3}\}$ (12)
    where $\mathbf{y}_{i}$ and $\mathbf{\hat{y}}_{i}$ are ground truth and prediction of lab results and $i$ means the $i$-th lab category.
  • Individual Training. We individually update parameters in decoder $\Theta_{d}$ after maintaining $\Theta_{e}$ unchanged given from the former phase, where we get refined lab embeddings $\mathbf{\tilde{l}_{1}}$, $\mathbf{\tilde{l}_{2}}$, and $\mathbf{\tilde{l}_{3}}$ through updated decoder.
    $L_{i}=\text{BCE}(\mathbf{\hat{y}}_{i},\mathbf{y}_{i}),\>i=1,2,3$ (13)
After two training phases, we get pretrained parameter $\Theta=[\Theta_{d};\Theta_{e}]$ for the predictive model. Similar to the real scenario, doctors always provide diagnoses relying on their historical conditions and abnormal lab results. By training model to understand such underlying laboratory pattern, it provides a pathway to deeply integrate additional laboratory knowledge into the model, ensuring that the predictions are not only accurate but also clinically meaningful.
3.5 Fine-Tuning and Inference
The proposed model can be trained with EHR data either with or without lab tests, making it adaptable to data missing desired features. If training with only diagnostic information, we can directly compute the estimated output $\hat{\mathbf{y}}^{\prime}$ and the downstream loss $\mathcal{L}^{\prime}$ for optimization:
$\hat{\mathbf{y}}^{\prime}_{\text{direct}}$ $=\sigma(\mathbf{W}\mathbf{p})\in\mathbb{R}^{o}$ (14)
$L^{\prime}_{\text{direct}}$ $=\text{Loss}(\mathbf{y}^{\prime},\hat{\mathbf{y}}^{\prime})$ (15)
If lab features are included, the model leverages pretrained parameters $\Theta$ to get both adjusted patient embedding $\tilde{\mathbf{p}}$ and lab embeddings $\mathbf{\tilde{l}_{1}}$, $\mathbf{\tilde{l}_{2}}$, and $\mathbf{\tilde{l}_{3}}$. The downstream outputs $\hat{\mathbf{y}}^{\prime}$ and fine-tuning loss $\mathcal{L}^{\prime}$ are then calculated as follows:
$\hat{\mathbf{y}}^{\prime}_{\text{finetune}}$ $=\sigma(\mathbf{W}[\mathbf{\tilde{l}_{1}}:\mathbf{\tilde{l}_{2}}:\mathbf{\tilde{l}_{3}}]|\tilde{\mathbf{p}})\in\mathbb{R}^{o}$ (16)
$L^{\prime}_{\text{finetune}}$ $=\text{Loss}(\mathbf{y}^{\prime},\hat{\mathbf{y}}^{\prime}|\Theta)$ (17)
Here, $\mathbf{W}\in\mathbb{R}^{o\times p}$ is the weight matrix and [:]delimited-[]:[:][ : ] means concatenation. Note that $\mathbf{y}^{\prime}$, $\sigma$, $o$, and $\mathcal{L}^{\prime}$ depend on the specific tasks. During optimization with backpropagation, the embedding matrix and parameters remain learnable.
4 Experiments
4.1 Experiment Setting
4.1.1 Datasets
To evaluate our proposed model, we focus on two pubic and widely-used EHR datasets: MIMIC-III Johnson et al. (2016) and MIMIC-IV Johnson et al. (2020). We focus on MIMIC-III as the primary dataset for our experiment and divide it into 6000/500/1000 for the training, validation, and test sets, respectively. The basic statistics of MIMIC-III and MIMIC-IV and results of MIMIC-IV are shown as supplement at the Appendix E.
Note that, we selected MIMIC-III as the primary dataset for two key reasons: it facilitates direct comparison with existing baselines commonly evaluated on MIMIC-III, ensuring reproducibility, and it natively supports ICD-9 codes, aligning with knowledge encoding base in baselines and our Diagnosis KG. Using MIMIC-IV would require mapping from ICD-10, potentially introducing errors that could affect model evaluation.
4.1.2 Baselines
To evaluate the performance of our proposed model, we selected the following 10 EHR models as comparison methods: (1) RNN/CNN-based models: Dipole Ma et al. (2017), Deepr Nguyen et al. (2017), RETAIN Choi et al. (2016a) and Timeline Bai et al. (2018). (2) Graph-based models: GRAM Choi et al. (2017), KAME Ma et al. (2018), and CGL Lu et al. (2021). (3) Transformer-based models: G-BERT Shang et al. (2019), HiTANet Luo et al. (2020), and THAM Li et al. (2024).
Note that although GraphCare Jiang et al. (2023) is another KG-based method, it is not included as a baseline in experiments because its KGs cannot be directly transferred from CCS codes to ICD-9 codes due to limited GPU memory. Supportive details will be provided in the following case study. Implementation details are discussed in Appendix D.
4.1.3 Tasks & Evaluation Metrics.
We conduct our experiments on two tasks as downstream examples for clinical prediction, following the similar settings as in previous studies Choi et al. (2017); Lu et al. (2021):
  1. 1.
    Diagnosis Prediction. This task involves multi-label prediction, where the goal is to predict all condition codes for the next admission based on admission history.
  2. 2.
    Heart Failure Prediction (HF). This is a binary classification task that predicts whether patients will be diagnosed with heart failure in their next admission.
Given the label imbalance in EHR data, we use the weighted $F_{1}$ score (w-$F_{1}$) and recall at k ($R@k$) as evaluation metrics for diagnosis prediction and use the area under the ROC curve (AUC) and the F1 score for HF prediction.
4.2 Prediction Results
Table 2 demonstrates that DualMAR consistently outperforms existing baselines across both diagnosis and HF prediction tasks for MIMIC-III dataset. The proposed framework consistently performs the best, underlining the efficacy of external priors from clinical KGs for prediction.
Table 2: Diagnosis and Heart Failure Prediction.
Diagnosis Prediction HF Prediction
Models w-$\mathbf{F_{1}}$ R@10 R@20 AUC $\mathbf{F_{1}}$
Dipole 14.66(0.2) 28.73(0.2) 29.44(0.2) 82.08(0.3) 70.35(0.2)
Deepr 11.68(0.1) 26.47(0.1) 27.53(0.1) 81.36(0.1) 69.54(0.1)
RETAIN 18.37(0.8) 32.12(0.8) 32.54(0.6) 83.21(0.3) 71.32(0.2)
Timeline 20.46(0.2) 30.73(0.1) 34.83(0.1) 82.34(0.3) 71.03(0.2)
GRAM 20.78(0.1) 34.17(0.2) 35.46(0.3) 81.55(0.2) 68.78(0.1)
KAME 21.10(0.1) 29.97(0.2) 33.99(0.2) 82.88(0.1) 72.03(0.1)
CGL 22.63(0.2) 33.64(0.3) 37.87(0.2) 84.19(0.2) 71.77(0.1)
G-BERT 22.28(0.3) 35.62(0.2) 36.46(0.2) 81.50(0.2) 71.18(0.1)
HiTANet 23.15(0.2) 34.68(0.3) 35.97(0.1) 85.13(0.1) 73.15(0.2)
THAM 23.92(0.1) 38.64(0.2) 39.97(0.1) 83.24(0.1) 74.12(0.1)
DualMAR 25.37${}_{(\textbf{0.3})}$ 40.52${}_{(\textbf{0.2})}$ 41.86${}_{(\textbf{0.3})}$ 86.53${}_{(\textbf{0.1})}$ 75.35${}_{(\textbf{0.1})}$
Table 3: Real-time and Non-chronic Diagnosis Prediction.
RT Diagnosis NC Diagnosis
Models w-$\mathbf{F_{1}}$ R@10 R@20 w-$\mathbf{F_{1}}$ R@10
Dipole 18.45(0.2) 29.87(0.2) 31.23(0.3) 13.83(0.7) 24.12(0.5)
Deepr 18.78(0.2) 30.45(0.2) 32.14(0.3) 10.56(0.9) 22.74(0.3)
RETAIN 19.56(0.2) 32.18(0.3) 33.67(0.3) 18.22(0.5) 30.15(0.4)
Timeline 21.12(0.2) 33.05(0.3) 34.23(0.3) 18.04(0.3) 29.58(0.4)
GRAM 22.54(0.2) 35.67(0.3) 37.10(0.4) 18.95(0.3) 30.08(0.3)
KAME 23.21(0.1) 36.51(0.1) 37.75(0.3) 19.13(0.2) 30.34(0.4)
CGL 24.13(0.1) 38.21(0.3) 39.64(0.4) 19.45(0.3) 30.52(0.5)
G-BERT 24.35(0.3) 37.12(0.5) 39.35(0.4) 19.26(0.2) 31.96(0.3)
HiTANet 25.67(0.3) 39.72(0.4) 41.85(0.4) 18.81(0.3) 31.08(0.4)
THAM 24.52(0.4) 40.31(0.2) 41.45(0.3) 19.93(0.3) 33.53(0.4)
DualMAR 26.89${}_{(\textbf{0.3})}$ 42.86${}_{(\textbf{0.1})}$ 43.15${}_{(\textbf{0.2})}$ 21.69${}_{(\textbf{0.2})}$ 34.35${}_{(\textbf{0.4})}$
Prediction Results on MIMIC-III. We report the average performance (%) and standard deviation (in bracket) of each model over 10 runs. The best results for each metrics are highlighted.
4.2.1 Real-time & Non-chronic Diagnosis Prediction.
Table 3 shows the results of two supplementary experiments, real-time and non-chronic diagnosis predictions, respectively.
  1. 1.
    Real-time (RT) Diagnosis: To fairly compare performance of all baselines, we feed lab tests information for all baselines and proposed model, which not only augments representation for diagnosis prediction but also mimics real-time diagnosis of physician within the beginning of an admission. We can observe that the proposed model still outperforms all baselines, which has even a larger improvement on w-F1 over regular diagnosis. We suppose that DualMAR, which originally relies on lab tests for auxiliary diagnosis, will be more suitable for lab features input than other “uni-aspect” baselines.
  2. 2.
    Non-chronic (NC) Diagnosis: Since MIMIC-III is mostly ICU data, we remove all chronic diseases in predictions and focus on predicting acute condition codes for future visits. We can observe that DualMAR still achieves superior performance, but all baselines and proposed model get relatively low values of metrics compared to regular diagnosis. Note that, R@20 will be omitted in report since the average number of diagnoses is less than 10 within each admission. We conjecture that the removal of chronic diseases in admissions skews the distribution, making rare diseases more prominent and, consequently, making the prediction task more challenging.
4.3 Ablation Study
To evaluate the effectiveness of involved modules in DualMAR, we compare performance of variants with some modules removed or replaced. We use both future diagnosis and HF prediction tasks on MIMIC-III as examples:
  1. 1.
    DualMARa: Removing Diagnosis KG and KG embedding modules, initializing embeddings randomly.
  2. 2.
    DualMARb: Replacing the embedding of the Diagnosis KG with GloVe, as used in GRAM Choi et al. (2017), to only focus on hierarchical information.
  3. 3.
    DualMARc: Replacing the embedding of the Diagnosis KG with text-embedding-3-large OpenAI (2024) to only retrieve semantic information.
  4. 4.
    DualMARd: Removing the Graph Learning module and directly feeding KG embeddings into the Attention layer.
  5. 5.
    DualMARe: Replacing EHR graph by a simplified graph with disease complications, excluding lab test nodes.
  6. 6.
    DualMARf: Replacing Attention by GRU modules.
  7. 7.
    DualMARg: Removing the proxy-task learning module and training the model directly on the downstream task.
Table 4: Ablation Results of the Variants of DualMAR.
Diagnosis Prediction HF Prediction
Model w-F1 R@10 R@20 AUC F1
DualMARa 21.21(0.2) 35.25(0.3) 36.46(0.3) 80.75(0.4) 69.81(0.3)
DualMARb 22.32(0.3) 36.28(0.3) 36.95(0.4) 81.66(0.3) 70.01(0.3)
DualMARc 23.15(0.3) 37.78(0.4) 38.79(0.4) 83.16(0.3) 72.73(0.3)
DualMARd 22.49(0.3) 34.86(0.3) 35.79(0.4) 80.90(0.3) 69.24(0.3)
DualMARe 24.58(0.3) 39.78(0.4) 40.46(0.4) 83.92(0.3) 73.62(0.3)
DualMARf 24.96(0.3) 40.16(0.3) 40.87(0.4) 85.81(0.3) 73.96(0.3)
DualMARg 23.74(0.3) 38.19(0.3) 39.52(0.4) 84.73(0.3) 72.85(0.3)
DualMAR 25.37${}_{(\textbf{0.3})}$ 40.52${}_{(\textbf{0.3})}$ 41.86${}_{(\textbf{0.3})}$ 86.53${}_{(\textbf{0.3})}$ 75.35${}_{(\textbf{0.3})}$
- no lab 24.21(0.3) 39.47(0.3) 40.35(0.3) 85.73(0.3) 73.92(0.3)
We report the results of DualMAR and its variants DualMARa-g in Table 4. The full DualMAR model consistently outperforms all variants, proving the effectiveness of our framework The comparisons between DualMAR and DualMARb-c highlight the importance of both hierarchical and semantic information from the KG, even though each alone can enhance predictions based on DualMARa. Additionally, the graph learning DualMARd and attention modules DualMARf are critical for refining representations, with the inclusion of lab tests further enhancing model performance in terms of DualMARe. The significant improvement of DualMAR over DualMARg underscores the essential role of the proxy-task learning module.
Notably, for cases without lab test data, the No LAB model, which uses modified GNN and skips pretraining, still achieves top-tier performance among baselines, demonstrating the effectiveness and generalizability of DualMAR.
4.4 Case Study
4.4.1 Less-frequent Diagnosis Prediction
We also design experiments focusing on predicting less frequent condition codes in MIMIC-III, similar to rare ICD code prediction tasks. We select and predict 100 ICD codes that occur fewer than 10 times in the dataset. We follow the previous settings which excludes records without rare codes to stabilize training. We use recall at 5/8/15/20 for evaluation. Figure 2(a) compares DualMAR with baselines using ontology or transformer architectures. DualMAR outperforms other selected methods on R@5, R@8, R@15, and R@20. Notably, CGL and G-BERT show significant improvements, indicating that external knowledge bases can aid in diagnosing rare conditions. This underscores the importance of incorporating medical ontologies or complex KGs to enhance model resilience against EHR data limitations.
(a) R@k in Less-frequent Diagnosis Prediction.
(b) w-$F_{1}$ of Diagnosis Prediction with different KGs.
Figure 2: Results for the $1^{st}$ and $2^{nd}$ case studies, full results are shown in Appendix F.
4.4.2 Advanced Analysis of KG Embeddings
To assess the importance of each component in our Diagnosis KG, we divided it into three parts: ICD9-Hierarchy KGa, Ontology-KG KGb, and LLM-KG KGc. Note that, the reason why both KGa+b and KGa+b+c are ignored in experiment is that knowledge in KGa is mostly covered by KGb, so that they will get same results as other existing settings. We then evaluated the predictive performance of the model with these varying complexities. Figure 2(b) shows how the complexity and embedding dimensions of KGs influence diagnosis prediction. Key observations include: (1) LLM-generated triples (KGc) enhance comprehensive KG representation, as seen in KGa+c and KGb+c. (2) Ontology-KG (KGb) from diverse data sources is more reliable for accurate predictions than ICD9-Hierarchy (KGa). (3) Embedding dimensions significantly impact predictions; higher dimensions offer a deeper understanding but can lead to overfitting, especially with simpler KGs like KGa.
4.4.3 Accessibility
We evaluated the accessibility of DualMAR with large-scale KGs, comparing it to GraphCare Jiang et al. (2023) using an NVIDIA RTX-4090 GPU with 24GB of memory. For a fair comparison, we adapted UMLS-GPT-KG in GraphCare to our Diagnosis KG, adjusting only the batch size due to its end-to-end structure without pretraining module. Table 5 shows average GPU usage for diagnosis prediction. GraphCare struggles with training on common commercial GPUs when transitioning from CCS KG to ICD-9 KG, even with small batch sizes, limiting its scalability with large KGs. In contrast, DualMAR efficiently separates learning processes for KGs and EHR data, allowing it to scale to larger KGs and making it more deployable across various EHR tasks and datasets.
Table 5: Accessibility Analysis with Adjusted GraphCare.
Models GPU Memory Usage Runnable
DualMAR - batch sizes: 256, 32, and 32
     - KG Embedding 11.73 GB
     - Pretrain & Finetune 9.78 GB
     - Direct Training 9.62 GB
GraphCare + Diagnosis KG
     - Batch Size: 2 33.95 GB ×
     - Batch Size: 4 42.57 GB ×
     - Batch Size: 8 98.23 GB ×
     - Batch Size: 16 172.04 GB ×
5 Conclusion
In this study, we introduced DualMAR, a framework designed to enhance EHR prediction tasks by leveraging both individual health records and public knowledge bases. Through the construction of a bi-hierarchical diagnosis KG augmented by LLMs and the design of a proxy-task learning via lab results, DualMAR significantly improves the representation of patient embeddings and disease-related predictions. Our experiments demonstrate that DualMAR outperforms existing state-of-the-art models, highlighting the potential of integrating comprehensive hierarchical and semantic information from KGs into EHR models. This approach not only broadens the scope of medical-domain knowledge but also addresses the limitations of traditional EHR models by incorporating more complex and nuanced medical data. Future work could explore expanding DualMAR to adapt to large-scale KGs and apply it to a broader range of healthcare prediction tasks beyond diagnostics. Additionally, this work highlights the crucial role of laboratory data in clinical decision-making, suggesting that further investigation into leveraging lab tests as auxiliary inputs for both diagnosis and treatment predictions remains a promising direction.
Appendix
Appendix A KG Construction
A.1 Data Sources in Healthcare KG.
To develop a comprehensive KG for studying diseases, we consider two primary resources related to diseases and phenotypes, and incorporate three additional resources to enhance the connectivity of our Diagnosis KG. Notably, these data resources—whether widely-used standardized ontologies or direct readouts from experimental measurements—primarily focus on extensive coverage across disease, phenotype, and drug entities. We begin with a brief introduction of primary data resources used to construct component of the Diagnosis KG:
  1. 1.
    DrugBank. DrugBank Knox et al. [2024] is a comprehensive resource containing detailed pharmaceutical knowledge. We retrieved the latest version (5.1.12), published on March 14, 2024. Our focus is on synergistic drug interactions, which represent the bidirectional connections between two drugs.
  2. 2.
    DrugCentral. DrugCentral Ursu et al. [2016] is a curated database that provides information on drug-disease interactions, including indications, contraindications, and off-label uses. We utilized the updated SQL Database, released on November 1, 2023, for our study.
  3. 3.
    HPO. The Human Phenotype Ontology Gargano et al. [2024] offers detailed information on phenotype abnormalities associated with diseases. We used one of the most recent updates, from April 19, 2024, focusing on disease-phenotype and phenotype-phenotype relationships, and extracted verified associations with expertly curated annotations.
  4. 4.
    ICD-9-CM Disease Ontology. ICD-9-CM Organization et al. [1988] is a coding system used by health insurers to classify medical conditions for billing purposes, and it underlies the diagnostic information in most EHR data. For our purposes, we extracted the parent-child relationships among codes to describe disease interactions.
  5. 5.
    SIDER. The Side Effect Resource Kuhn et al. [2016] contains data on side-effect phenotype caused by various drugs. We retrieved this information from the data release dated October 21, 2015.
A.2 Data Sources in LLM-generated KG.
We also extract textual features for disease nodes in the KG from several widely-used clinical knowledge bases. The in-context learning capabilities of LLMs are employed to convert this textual information into triples, which are then integrated into the KG. The following four public data sources are used for triple generation:
  1. 1.
    Mayo Clinic. Mayo Clinic Brennan et al. [1998] is a nonprofit academic medical center and biomedical research institution that maintains a knowledge base with clinical information on over 2000 diseases. We collected Mayo Clinic web knowledge in 2023 and mainly focus on both “Symptoms & Causes” (Overview, Symptoms, When to see a doctor, Causes, Risk Factors, Complications, and Prevention) and “Diagnosis & Treatment” parts for each condition.
  2. 2.
    Orphanet. Orphanet Weinreich et al. [2008] is a database dedicated to rare diseases. We collected data in 2023, focusing on information about definitions, prevalence, treatment, and clinical descriptions for rare diseases.
  3. 3.
    Rare Disease Database. Rare Disease Database Schieppati et al. [2008], created by a nonprofit organization NORD, serves as a supplementary source for rare diseases. In 2023, we collected data to retrieve information on disease overviews, symptoms, causes, treatments, and clinical trials.
  4. 4.
    Wikipedia. Wikipedia is a free, open-content online encyclopedia maintained by volunteers worldwide. It includes an official list of ICD-9 condition codes. We web-scraped the overall hierarchy of ICD-9 codes and gathered related information for diseases with individual Wikipedia pages through valid hyperlinks. For example, the page for “Typhoid Fever” provides detailed information on symptoms, causes, diagnosis, prevention, treatment, epidemiology, history, terminology, and societal impact, offering more comprehensive and diverse knowledge compared to other textual sources.
Table 6: General Prompting Framework for Triple Generation Tasks of LLMs
Name Prompt Template
Variables <category>: The prompt is not only suitable for disease but also available for concepts within medical domains such as medication and treatment. In this study, only condition concepts are considered. <term>: The concept name. In this study, it means disease names provided by web-scraped text. <topics>: The topics related to crawled text, and it has been provided when we crawled them from websites. <text>: The content of crawled text.
Skeleton Given a crawled text about specific topic of certain <category>, please find triples related to the given <category> in terms of crawled text. Filling triples in updates based on given information and strictly following output style of example updates. Each update should follow the format of [ENTITY 1, RELATIONSHIP, ENTITY 2] with directed edge. Both ENTITY 1 and ENTITY 2 should be noun, and one of them must be <term>. Just output each unique triple once, don’t output repeatedly. It is possible that <category> name not exactly matched in crawled text (abbreviated or partly matched), consider it as the same thing. Example: ## An example demo is shown below… Given a paragraph about specific topic of certain <category>, please find triples related to the given <category> in the text. Given Information: <category> Name: <term> Topics: <topics> Text: <text> Updates: ## LLM’s output is expected to generate here in terms of given information… ## [Head Entity, Relation, Tail Entity]
Example Disease Name: Heart Failure Topics: Overview Text: Heart failure occurs when the heart muscle doesn’t pump blood as well as it should. When this happens, blood often backs up and fluid can build up in the lungs, causing shortness of breath. Certain heart conditions gradually leave the heart too weak or stiff to fill and pump blood properly. These conditions include narrowed arteries in the heart and high blood pressure. Proper treatment may improve the symptoms of heart failure and may help some people live longer. Lifestyle changes can improve quality of life. Try to lose weight, exercise, use less salt and manage stress. But heart failure can be life-threatening. People with heart failure may have severe symptoms. Some may need a heart transplant or a device to help the heart pump blood. Heart failure is sometimes called congestive heart failure. Updates: [Heart Failure, IS_CAUSED_BY, Narrowed Arteries], [Heart Failure, IS_CAUSED_BY, High Blood Pressure], [Heart Failure, HAS_SYMPTOMS, Shortness of Breath], [Heart Failure, HAS_SYMPTOMS, Fluid Build-up in Lungs], [Heart Failure, NEEDS_TREATMENT, Proper Treatment], [Heart Failure, NEEDS_TREATMENT, Lifestyle Changes] ## You can use more examples to refine output of LLMs, but only one example is also fine for this task.
A.3 Prompting LLM
It is important to note that since all textual information for disease nodes is sourced from either publicly verified or expertly curated data sources, issues of data quality or hallucination in the medical knowledge generated by the LLM are not a primary concern. In this section, we focus on how to utilize the in-context learning capabilities of LLMs to transform this textual information into structured triples that can be directly integrated into the KG.
The LLM used in this study is Llama-3-8b. Table 6 illustrates a carefully designed prompt for constructing triples from disease conditions. The primary objective is to harness the extensive knowledge embedded in the LLM to extract meaningful triples consisting of two entities and a relationship. Llama-3-8b was chosen for inference in this study for several reasons: first, its relatively small parameter size makes it easier to deploy and use; second, it is one of the latest and most widely adopted open-source models; and third, the task at hand is common and straightforward, not requiring the large parameter size or fine-tuning that other medical-domain LLMs might offer.
Our strategy begins with a prompt related to a specific medical condition. The LLM is then tasked with generating a list of triples, following the format [ENTITY 1, RELATIONSHIP, ENTITY 2], where both ENTITY 1 and ENTITY 2 are nouns. The goal is to generate these triples with both breadth (capturing potential abbreviations or partial matches) and accuracy (ensuring that each triple is based on existing information).
Although triple generation is relatively straightforward for LLMs, it still remains two issues:
  • Sometimes, the output is either empty or consists of garbled text, which we consider as deprecated results (approximately 8% of cases).
  • Occasionally, the output triples are incomplete or end with “…” (approximately 3% of cases).
To mitigate these issues, we employ iterated running and re-read prompting strategies to ensure stable outputs. For iterated running, we set the hyperparameter $x$ to control the number of times the entire textual data is re-prompted, with $x=2$ in this study. For re-read prompting, we set the hyperparameter $y$ to repeatedly feed the same prompt into the model, allowing the LLM to reconsider and refine its answers, with $y=1$ in this study. These strategies help the LLM generate comprehensive triples from all textual information, with the final output curated through manual checks.
Considering accessibility, we chose to utilize open-source LLMs for the triple generation step, as they are more accessible and convenient to deploy on local servers compared to the GPT series. In this experiment, while Llama-3-8b serves as our inference LLM example, other widely-used LLMs like Llama-2-13b or fine-tuned models like PMC-Llama can also be used in this process.
Appendix B KG Embeddings Comparison
Most current hierarchical KG embeddings are designed to capture and represent hierarchy information in KGs using various geometric structures, including Euclidean and non-Euclidean geometries. Euclidean-based methods typically focus on projections across different coordinates to distinguish entities at various levels. Compared to non-Euclidean approaches like hyperbolic and spherical embeddings, Euclidean projections can efficiently retrieve hierarchical information without additional processes or high computational costs. Additionally, we can separately train embeddings through link prediction tasks, allowing GPU usage to be decoupled from the graph learning process, ensuring stable performance even with more complex KGs. We propose projecting each entity in the KG onto polar coordinates, which enables more efficient embedding retrieval.
Table 7 compares the polar-space-based KG embedding method, which is an adjusted embedding method refered from HAKE Zhang et al. [2020], with five widely-used baselines Bordes et al. [2013], Sun et al. [2019], Nickel and Kiela [2017], Fionda and Pirrò [2020], demonstrating the advantages of our approach in retrieving both hierarchical and semantic features from the KG. Performance is evaluated using MRR and HITS@k metrics for regular link prediction tasks, with k set to 1, 3, and 10 in our experiments.
Table 7: Different Embedding Baselines Comparison (%)
MODEL MRR HITS@1 HITS@3 HITS@10
TransE 54.77 45.90 58.30 72.40
RotatE 57.82 47.95 62.56 75.28
ModE 63.02 55.46 66.25 77.85
TripleRE 64.02 57.21 68.03 79.32
HyperE 59.03 53.69 64.82 74.93
HAKE 67.37 60.23 70.62 81.21
Appendix C Details in Multilevel Attention
After generating $\mathbf{X}$ for all medical codes appears in the intersection of both training set $\mathcal{S}$ and concept vocabulary $\mathcal{C}$, our objective is to utilize EHR dataset to train both the model parameters and the embedding matrix $\mathbf{E}\in\mathbb{R}^{|\mathcal{C}|\times d}$ for medical codes, and $d$ is the pre-defined embedding dimension for the previous GNN module. For the training process, we only focus on each patient $u$ with multiple admission records. An encoder is employed to transform all of their admission records into a patient representation vector $\mathbf{p}$, leveraging the embeddings in $\mathbf{X}$:
$\mathbf{p}=\text{Bi-Attention}(\mathbf{V}_{1},\mathbf{V}_{2},\ldots,\mathbf{V}_{T}|\mathbf{X}).$
In this study, “Bi-Attention” means that we implement a two-tier attention mechanism, detailed as follows:
a) Learning Admission Embeddings: Consider an admission record $\mathbf{V}_{\tau}$ containing $n$ medical codes. The embedding $\mathbf{x}_{i}$ for each medical code $c_{i}\in\mathbf{V}_{\tau}$ can be retrieved using $\mathbf{X}$, so that we can get code embeddings $\mathbf{E}^{\tau}_{V}$, which is also the subset of embeddings $\mathbf{E}$, for each admission. To aggregate these code embeddings into an admission embedding $\mathbf{v}_{\tau}$, we apply a global attention mechanism:
$\mathbf{E}^{\tau}_{V}=[\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{n}]\in\mathbb{R}^{n\times m}$
$\mathbf{z}_{i}=\tanh(\mathbf{W}_{c}\mathbf{x}_{i})\in\mathbb{R}^{a}$
$\mathbf{v}_{\tau}=\sum_{i=1}^{n}\frac{\exp(\mathbf{z}_{i})}{\sum_{j=1}^{n}\exp(\mathbf{z}_{j})}\mathbf{x}_{i}\in\mathbb{R}^{m}.$
Here, $\mathbf{W}_{c}\in\mathbb{R}^{a\times m}$ is the weight matrix, where $a$ represent the attention dimension. The attention scores $\alpha{\tau}=[\alpha_{\tau}^{1},\alpha_{\tau}^{2},\ldots,\alpha_{\tau}^{n}]$ quantify the contribution of each medical code within an admission, and the weighted sum of $\mathbf{x}_{i}$ results in the admission embedding $\mathbf{v}_{\tau}$.
b) Learning Patient Embeddings: Once the embedding $\mathbf{v}_{\tau}$ for the $\tau$-th admission is computed, we derive the patient embedding by aggregating across all admissions. The admission embedding $\mathbf{v}_{\tau}$ is first projected into the patient embedding space through a simple linear classifier:
$\tilde{\mathbf{v}}_{\tau}=\sigma(\mathbf{W}_{u}\mathbf{v}_{\tau})\in\mathbb{R}^{p}.$
Here, $\mathbf{W}_{u}\in\mathbb{R}^{p\times m}$ is the corresponding weight matrix, and $p$ denotes the dimension of the patient embedding. The activation function is denoted as $\sigma$.
Assume patient $u$ has $T$ admission records as input features of model, we can get the admission embeddings $\mathbf{E}^{u}_{\mathbf{p}}$ after the code-level attention module. To aggregate $\mathbf{E}^{u}_{\mathbf{p}}$ into the final patient embedding $\mathbf{p}$, we then also apply the global attention across different admissions for each patient:
$\mathbf{E}^{u}_{\mathbf{p}}=[\tilde{\mathbf{v}}_{1},\tilde{\mathbf{v}}_{2},\dots,\tilde{\mathbf{v}}_{T}]\in\mathbb{R}^{T\times p}$
$\mathbf{r}_{\tau}=\tanh(\mathbf{W}_{v}\tilde{\mathbf{v}}_{\tau})\in\mathbb{R}^{b}$
$\mathbf{p}=\sum_{\tau=1}^{T}\frac{\exp(\mathbf{r}_{\tau})}{\sum_{\tau=1}^{T}\exp(\mathbf{r}_{\tau})}\tilde{\mathbf{v}}_{\tau}\in\mathbb{R}^{p}.$
In this case, $\mathbf{W}_{v}\in\mathbb{R}^{b\times p}$ is the weight matrix used for the transformation. The attention scores $\beta=[\beta_{1},\beta_{2},\ldots,\beta_{T}]$ capture the relative importance of each admission, and their weighted sum produces the final patient embedding $\mathbf{p}$.
Table 8: Statistics of MIMIC-III and MIMIC-IV datasets
Dataset MIMIC-III MIMIC-IV
# patients 7,493 10,000
Max. # visit 42 55
Avg. # visit 2.66 3.66
# codes 4,880 6,102
Max. # codes per visit 39 50
Avg. # codes per visit 13.06 13.38
Appendix D Experiment Implementation
D.1 Data Processing Details
To evaluate our proposed model, we focused on two well-established datasets in critical care research: MIMIC-III and MIMIC-IV. Both datasets are derived from de-identified clinical data collected at the Beth Israel Deaconess Medical Center in Boston, Massachusetts, and include detailed records from patients admitted to the Intensive Care Units (ICUs). MIMIC-III contains data from over 40,000 ICU admissions between 2001 and 2012, covering a wide range of information such as patient demographics, vital signs, laboratory results, diagnoses, and diagnostic codes. MIMIC-IV extends this dataset to approximately 60,000 ICU admissions from 2008 to 2019, reflecting more recent clinical practices and patient demographics.
To ensure distinct patient cohorts, we selected patients from MIMIC-IV who were admitted between 2018 and 2019, avoiding temporal overlap with MIMIC-III. We included only patients with multiple visits (# of visits $\geq$ 2) to ensure sufficient data for analysis. Specifically, for MIMIC-III, we divided the data into 6000 training, 500 validation, and 1000 testing samples. For MIMIC-IV, the distribution comprised 8000 training, 1000 validation, and 2000 testing samples. This approach allows us to leverage the longitudinal structure of EHR data, with the last visit of a patient serving as the label and all preceding visits as features. We aim to enhance the model’s ability to accurately predict critical care outcomes, contributing to advancements in predictive modeling and ultimately improving patient care through data-driven insights.
D.2 Parameter Settings
We consider node feature of graph learning module as the hierarchy KG embedding upon polar-space, which is $20002k=$ 2000 dimension vectors. Specifically, the embedding size $k=1000$ for both hierarchical and semantic representation in Diagnosis KG. We use two Graph Attention (GAT) Convolution layers for disease-lab graph, where the hidden dimension for both layers are 256, and we remain other settings as default values. The attention sizes $a=256$ and $b=256$ remain the same dimension as previous output, and to avoid over-fitting problem, we set dropout layers within both attention layers with $0.2$ dropout rate. Moreover, three three-level MLPs are adopted as decoders to help us get lab test embeddings with uniformly 256, 128 hidden dimensions, and we set dropout rate between layers are all 0.4 for either pretraining or direct training process. In addition, to finetune the lab test embeddings for various downstream tasks, a two-level classifiers with 256 hidden dimension and 0.5 dropout rate are used for transforming concatenated features to logits.
To retrieve precise hierarchical and semantic information from Diagnosis KG, we modify the source code of HAKE Zhang et al. [2020] and set training step as 180,000. For proxy-task learning, we use 10 and 10 epochs for joint and individual training respectively. We uniformly use Adam optimizer and decay learning rate for all training process, and we use 500 and 50 epochs for diagnosis prediction and heart failure prediction, respectively. Moreover, the initial learning rate in both pretraining and finetuning processes is 0.001. All programs are implemented using Python 3.10 and Pytorch 2.3.1 with CUDA 12.4 on a machine with two AMD EPYC 9254 24-Core Processors , 528GB RAM, and four Nvidia L40S GPUs.
Table 9: Prediction Results on MIMIC-IV for Diagnosis and HF Prediction.
Diagnosis Prediction HF Prediction
Models w-$\mathbf{F_{1}}$ R@10 R@20 AUC $\mathbf{F_{1}}$
Dipole 22.16(0.2) 36.21(0.2) 38.74(0.2) 84.80(0.3) 69.52(0.2)
Deepr 22.58(0.1) 36.79(0.1) 39.45(0.1) 83.61(0.5) 70.46(0.1)
RETAIN 23.11(0.8) 37.32(0.8) 40.15(0.6) 84.14(0.3) 71.23(0.2)
Timeline 23.76(0.2) 37.89(0.1) 40.87(0.1) 83.45(0.3) 72.30(0.2)
GRAM 24.39(0.1) 38.42(0.2) 41.62(0.3) 85.55(0.2) 69.82(0.4)
KAME 25.01(0.1) 38.86(0.2) 42.12(0.2) 84.80(0.2) 72.34(0.5)
CGL 25.74(0.2) 39.23(0.3) 42.67(0.2) 87.91(0.2) 70.71(0.3)
G-BERT 25.12(0.3) 39.91(0.2) 43.25(0.2) 85.76(0.2) 72.88(0.1)
HiTANet 24.53(0.2) 38.42(0.3) 41.89(0.1) 86.34(0.4) 71.35(0.2)
THAM 26.97(0.3) 43.07(0.4) 47.19(0.2) 87.43(0.3) 72.26(0.2)
DualMAR 29.87(0.3) 45.66(0.2) 51.73(0.3) 90.32(0.2) 73.54${}_{(\textbf{0.1})}$
Appendix E Details in Experiment
Table 8 shows the basic statistics for MIMIC-III and MIMIC-IV. Table 9 shows the result of main experiment which is the same setting as shown in MIMIC-III. Note that, we remain DualMAR in the same hyperparameter setting as model in MIMIC-III, but it still achieve superior performance across all baselines, which display the effectiveness and robustness of proposed model. Compared to MIMIC-III, MIMIC-IV has larger sample size and more longitude data, which accelerate the predictive performance of DualMAR within more complex scenarios, and we can tune the hyperparameters in model for pursuing even more accurate prediction for various downstream tasks.
Moreover, to evaluate the statistical significance of performance differences between our proposed DualMAR model and baseline models, we employed both the paired t-test and Wilcoxon signed-rank test. The t-test was chosen under the assumption of normally distributed performance metrics. To account for potential deviations from normality, the non-parametric Wilcoxon signed-rank test was also conducted. These tests allow us to robustly determine whether the observed improvements in DualMAR’s performance are statistically significant compared to the baselines.
Here we only take the w-$F_{1}$ of Diagnosis Prediction in MIMIC-IV as an example to show the evaluation process on Table 10. We observe that the Wilcoxon signed-rank test consistently yielded p-values of $1.95\times 10^{-3}$, indicating significant differences between DualMAR and all baseline models. The paired t-tests further confirmed these results, with all p-values well below $1\times 10^{-10}$. The 95% confidence intervals for the t-tests showed a clear positive difference, reinforcing the superior performance of DualMAR in Diagnosis Prediction compared to each baseline. These results demonstrate the robustness of DualMAR’s improvements in weighted $F_{1}$ scores over the baselines.
Table 10: Wilcoxon Signed-Rank Test and t-Test Results for DualMAR vs. Baseline Models with 10 runs. “p” means p-value for certain evaluation method, and “CI” means Confidence Interval calculated by T-testing method.
Models Wilcoxon p T-test p 95% CI
DualMAR vs. Dipole 1.95e-03 1.62e-14 (7.41, 7.80)
DualMAR vs. Deepr 1.95e-03 2.01e-14 (7.17, 7.53)
DualMAR vs. RETAIN 1.95e-03 6.27e-10 (6.46, 7.39)
DualMAR vs. Timeline 1.95e-03 9.77e-13 (5.93, 6.39)
DualMAR vs. GRAM 1.95e-03 5.53e-13 (5.31, 5.67)
DualMAR vs. KAME 1.95e-03 1.23e-12 (4.87, 5.21)
DualMAR vs. CGL 1.95e-03 3.46e-11 (4.49, 4.88)
DualMAR vs. G-BERT 1.95e-03 7.65e-13 (5.62, 6.01)
DualMAR vs. HiTANet 1.95e-03 1.99e-12 (5.89, 6.27)
DualMAR vs. THAM 1.95e-03 4.57e-11 (3.98, 4.32)
Table 11: Less Frequent Codes Diagnosis Prediction.
Rare-10 Rare-20
Model R@5 R@8 R@15 R@5 R@8 R@15
CGL 32.48 39.47 52.17 12.53 16.03 24.40
G-BERT 28.97 35.72 48.68 14.95 17.11 25.10
HiTANet 21.74 26.09 46.38 11.99 15.43 22.49
THAM 30.43 39.13 43.48 12.95 17.56 23.79
DualMAR 39.13 47.83 56.52 14.28 18.00 26.30
Table 12: Diagnosis Prediction Results by different KGs.
Models # Dims w-$F_{1}$ R@10 R@20
DualMAR-KGa 500 21.75(0.3) 33.27(0.4) 35.32(0.4)
DualMAR-KGa 2000 19.79(0.5) 32.12(0.4) 33.57(0.4)
DualMAR-KGb 500 23.87(0.3) 37.73(0.4) 39.18(0.4)
DualMAR-KGb 2000 24.58(0.3) 39.27(0.4) 40.32(0.4)
DualMAR-KGa+c 500 23.56(0.2) 37.43(0.3) 38.74(0.4)
DualMAR-KGa+c 2000 23.47(0.3) 37.89(0.4) 38.54(0.4)
DualMAR-KGb+c 500 24.32(0.3) 39.73(0.4) 40.15(0.4)
DualMAR-KGb+c 2000 25.37${}_{(\textbf{0.3})}$ 40.52${}_{(\textbf{0.2})}$ 41.86${}_{(\textbf{0.3})}$
Appendix F Details in Case Studies
We design experiments focusing on predicting less frequent condition codes in MIMIC-III, similar to rare ICD code prediction tasks. We conduct two subtasks:
  • Rare-20: Predicting 200 ICD codes that occur fewer than 20 times in MIMIC-III.
  • Rare-10: Predicting 100 ICD codes that occur fewer than 10 times in MIMIC-III.
Table 12 shows the full results of both Rare-10 and Rare-20. It highlights DualMAR’s superior performance in predicting less frequent ICD codes across both subtasks, Rare-10 and Rare-20. Specifically, DualMAR consistently outperforms baseline models, achieving the highest Recall at all thresholds (R@5, R@8, R@15). Notably, for Rare-20, DualMAR significantly surpasses other models, particularly in R@8 and R@15, reflecting its strong capability in handling rare condition codes. These results underscore the effectiveness of DualMAR in scenarios where capturing rare but critical diagnoses is crucial.
Table 12 shows the full results of DualMAR with varying KGs and embedding sizes. Beyond analysis for w-$F_{1}$ in the case study, full table also demonstrates the Recall performance of DualMAR. The results indicate that incorporating KG components (KGb+c) consistently yields the highest Recall across R@10 and R@20, particularly with 2000-dimensional embeddings. Notably, KGb also shows strong performance, suggesting the significant contributions of fundamental Ontology-KG in enhancing diagnostic predictions. The comparison highlights the critical role of augmenting KG to achieve superior recall in diagnosis prediction tasks.